# Index Terms: 5G Security, Wireless Intrusion Detection, Machine Learning, Principal Component Analysis, Software Defined Radio, Spectrograms.I Introduction

### v1

@@ Abstract
5G cellular networks are particularly vulnerable against narrowband jammers that target specific control subchannels in the radio signal. One mitigation approach is to detect such jamming attacks with an online observation system, based on machine learning. We propose to detect jamming at the physical layer with a pre-trained machine learning model that performs binary classification. Based on data from an experimental 5G network, we study the performance of different classification models. A convolutional neural network will be compared to support vector machines and -nearest neighbors, where the last two methods are combined with principal component analysis. The obtained results show substantial differences in terms of classification accuracy and computation time.
@@
5G Security, Wireless Intrusion Detection, Machine Learning, Principal Component Analysis, Software Defined Radio, Spectrograms.I Introduction
5G cellular networks promise improved latency, data rate, and the support of mission critical applications. Consequently, their application in automotive and smart manufacturing is rapidly increasing [1].
Unfortunately, 5G radio signals are particularly vulnerable to jamming attacks [2]. A jammer performs a denial of service attack by emitting interference in the frequency band used by, e.g., a 5G network. By targeting specific control subchannels in the signal, it is sufficient to interfere only with a small fraction of the overall bandwidth to disrupt the service. Such narrowband jammer can be easily implemented as an software defined radio (SDR) at low cost. One very effective narrowband jamming approach [3] targets the signal synchronization block (SSB) in the 5G signal, which is critical to establish and to maintain 5G communication. Other narrowband jamming approaches are discussed in [2].
One mitigation approach is to detect jamming attacks by an online wireless intrusion prevention systems (WIPS). Such detection often relies on specific parameters such as signal-to-noise-ratio (SNR), bit error rate (BER), packet error rate (PER) [4], or on measurements aggregated over multiple network layers [5]. Such high-level measurements, however, may be easily evaded by a jammer. For instance, a short jamming impulse in the SSB-subband effectively disrupts 5G communication [3] but barely affects the, usually time-averaged, SNR of the mainband. The same argument holds for SNR derivatives such as BER and PER. Detection based on measuring the quality of the radio signal at the physical layer [6], requires accurate synchronization in time and frequency. This not only leads to costly implementation but may also mislead the classification due to normal clock differences in a cellular network.
In this paper, we propose a WIPS to detect 5G narrowband jammers. This detection is based on spectrograms of the 5G radio signal and a pre-trained machine learning model for binary classification. These spectrograms are obtained from an experimental 5G network, which is occasionally attacked by an SSB jammer. The objective of our paper is to compare how accurate and how quickly the SSB jammer can be detected with different machine learning models. To this end, we compare classification accuracy and computational complexity of a convolutional neural network (CNN) with support vector machines (SVMs) and -nearest neighbors (KNNs). SVM and KNN are combined with principal component analysis (PCA) for dimensionality reduction.
@@
This network is monitored by a watchdog β a separate network element placed within range of the gNB. The watchdog receives radio signals and converts them to the digital baseband in the form of in-phase and quadrature (IQ) samples. Based on these IQ samples, anomaly detection is performed based on a pre-trained machine learning model.
@@
This configuration has several benefits. The watchdog can be deployed independently of the cellular network and, as a mere receiver, does not transmit signals that can be detected. Thus, it can be deployed invisible to a potential attacker. The watchdog is simpler and easier to construct than a communication device, since it can still operate effectively with imperfect synchronization in time and frequency.
II Dataset Description
We assume that the watchdog knows the basic radio parameters such as center frequency , bandwidth , and the pilot structure. This assumption is feasible, as these parameters are constant and known to the operator of the radio access network (RAN). Based on these parameters, the detector receives radio signals, computes a stream of IQ samples and spectrograms. This signal processing and the dataset creation are briefly explained in Sec. II-B.
@@ II Dataset Description
The first two cases are classified as legitimate and labelled with , while the third case is classified as an attack and labelled with .
II-A Experimental Setup
Fig. 1 shows the experimental setup. We are running a private 5G network in the n78 frequency band with center frequency βMHz. The system operates at 100βMHz bandwidth in time division duplexing (TDD) mode. The gNB implements the 5G new radio (NR) air interface using srsRAN 23.10 [7] and the universal software radio peripheral (USRP) n300 SDR frontend [8] connected to a personal computer (PC). The UE is a Quectel RM520N-GL modem [9], which is connected via USB 3.0 to a computer. The core network functionality is provided by Open5GS 2.6.6 [10], running on the same generic computer as srsRAN. This setting provides a 5G standalone network and complies with Release 17.4.0 of the 5G standard series 38 [11].
The jammer and the watchdog run on separate computers, each using one USRP X310 [12] RF frontend. The jammer permanently transmits uniform noise signal of over the SSB subchannel at βMHz with βMHz bandwidth. The watchdog permanently records IQ samples over βMHz bandwidth to include the spectrum of the neighboring bands.
Dataset creation, training and measuring the classification time was performed on a single workstation with an Intel Xeon w7-2495X CPU and an NVIDIA RTX A6000 GPU. The operation system Ubuntu 22.04 LTS with Python 3.9.12 and keras 2.12.0 allowed to use all CPU cores during classification. The GPU was only used for training and is, thus, not utilized during the measurement of classification time.
II-B Dataset Creation
A spectrogram is composed as a stack of PSDs arrays. Each PSD array is obtained by applying an FFT to a window of IQ samples. We used an FFT rather than Welchβs method [13], sacrificing precision for computational speed. A spectrogram is then composed as a matrix.
Based on previous measurement campaigns, we choose a window size of , leading to a frequency resolution of kHz and a stack of PSDs per spectrogram. This corresponds to a time window of ms.
The resulting spectrogram-matrices presented two problems when fed into the training set of the model. First, the power of received radio signals is very low in the linear domain and, thus, typically expressed in the logarithmic domain (decibel). Similarly, we apply the monotonic function to each value of the PSD array at each frequency, to avoid the vanishing gradient problem [14]. Second, due to approximation errors with very low energy bins, some values of the PSDs turned out to be , which causes computation errors with the log function. Then used a small constant , leading to the applied transformation . To each sample , a label is assigned.
@@ III-A Convolutional Neural Network
As spectrograms are two-dimensional, the first model of choice is a CNN whose structure is given in Table I. The objective is to return the correct label associated to each sample. The chosen loss function is the binary cross-entropy
@@
Defining as the class of no-jamming and as the class of jamming, the jammer detection is performed by the following test function on the input sample :
where is the chosen threshold. We, thus, measure accuracy as probability of false alarm (FA) and misdetection (MD) defined as
@@ III-B Principal Component Analysis
Using spectrograms as model input, leads to dimensions. To avoid the curse of dimensionality, we reduce the number of coordinates in two steps.
@@
Second, we further reduce the dimensions with principal component analysis (PCA). The basic idea of PCA is to change the basis so that the first coordinates (or principal components) contain βthe most informationβ about the underlying data. The ratio of explained variance measures the information contained in each component and sums up to , when summed over all components. The cumulated ratio of explained variance is displayed in Fig. 2. It can be seen that few components already capture most of the dataβs information. In particular, 98%, 98.5% and 99% of the variance are explained by , and principal components, respectively. This suggests components an optimal trade-off between computational complexity and accuracy.
@@ III-D Support Vector Machines
SVMs are a powerful class of machine learning models, which can be used for supervised classification and regression. SVMs operate by constructing a decision hyperplane in the feature space that optimally separates different classes. By employing the kernel trick, support vector machines can also be used effectively for non-linear tasks. In this work, we employed SVMs with varios kernels after dimension reduction with PCA.
@@ IV-A Convolutional Neural Network
The training set was composed of samples, equally distributed among the three cases described in Sec. II. Using the same distribution, the validation set was composed of samples. This set is used to monitor the validation loss and to stop the training phase after three increases of the parameter. The test set contains samples, distributed as the training and validation set.
@@
For each of the training, validation and test set, 1024 samples per class were selected from each of the three cases in Sec. II. The training, validation and test accuracy for the settings (i) standardized PSDs and (ii) based on projections of non-standardized PSDs onto the first 8 principal components are displayed in Table II. The optimal choice of neighbors based on the validation data is for setting (i) and for settings (ii). With an accuracy of approximately %, we can reliably detect jammers.
Fig. 5 shows the CDF of the classification time of the KNN. With standardization, the model classified of the cases within ms. With PCA, the model performed substantially better with an classification time of ms for of the cases.
@@ IV-C Support Vector Machines
Our experiments with SVMs showed that this class of models is well suited to detecting jammers. The overall accuracy on the validation and test dataset significantly depends on the used kernel. We considered linear, polynomial and radial basis function (RBF) kernels. On the other hand, increasing the number of dimensions does not show a significant effect.
@@
This large reduction of dimension, highly reduces the computational complexity of the classification. Fig. 6 shows that the SVM model classified 95% of the cases within ms. Thus, SVMs can detect jammers using substantially smaller computation time than the CNN (Fig. 4) and than the KNN without PCA (Fig. 5).
@@ V Conclusion
From an accuracy point of view the three models yielded similar results, showing that they could distinguish between the two classes with accuracy in the CNN case and with more than accuracy in the other two cases.
The results for computation time are interesting due to their substantial differences. While the CNN required 38 ms to classify 95% of the cases, KNN needed only 8 ms if combined with PCA. The fasted classification was achieved by SVM with PCA, with a classification time of ms for 95% of the cases.
Based on these promising results, we will now study the robustness of the detection accuracy for various wireless network geometries and for multiple jamming signals.

### v2

# Index Terms: 5G Security, Jamming, Wireless Intrusion Detection, Machine Learning, Principal Component Analysis, Software Defined Radio, Spectrograms.I Introduction

@@ Abstract
5G cellular networks are particularly vulnerable against narrowband jammers that target specific control subchannels in the radio signal. One mitigation approach is to detect such jamming attacks with an online observation system, based on machine learning. We propose to detect jamming at the physical layer with a pre-trained machine-learning model that performs binary classification. Based on data from an experimental 5G network, we study the performance of different classification models. A convolutional neural network will be compared to support vector machines and -nearest neighbors, where the last two methods are combined with principal component analysis. The obtained results show substantial differences in terms of classification accuracy and computation time.
@@
5G Security, Jamming, Wireless Intrusion Detection, Machine Learning, Principal Component Analysis, Software Defined Radio, Spectrograms.I Introduction
5G cellular networks promise improved latency, data rate, and the support of mission-critical applications. Consequently, their application in automotive and smart manufacturing is rapidly increasing [1].
Unfortunately, 5G radio signals are particularly vulnerable to jamming attacks [2]. A jammer performs a denial of service attack by emitting interference in the frequency band used by, e.g., a 5G network. By targeting specific control subchannels in the signal, it is sufficient to interfere only with a small fraction of the overall bandwidth to disrupt the service. Such narrowband jammer can be easily implemented as a software-defined radio (SDR) at low cost. One very effective narrowband jamming approach [3] targets the signal synchronization block (SSB) in the 5G signal [4], which is critical to establishing and maintaining 5G communication. Other narrowband jamming approaches are discussed in [2]; in particular in [5] it is documented how to locate in the frequency domain narrowband interference, but with no major effects on the reliability of the communication as in our case.
One mitigation approach is to detect jamming attacks by an online wireless intrusion prevention system (WIPS). Such detection often relies on specific parameters such as signal-to-noise-ratio (SNR), bit error rate (BER), packet error rate (PER) [6], orthogonality conditions[7], or on measurements aggregated over multiple network layers [8]. Such high-level measurements, however, may be easily evaded by a jammer. For instance, a short jamming impulse in the SSB-subband effectively disrupts 5G communication [3] but barely affects the, usually time-averaged, SNR of the main band. The same argument holds for SNR derivatives such as BER and PER. Detection based on measuring the quality of the radio signal at the physical layer [9], requires accurate synchronization in time and frequency. This not only leads to costly implementation but may also mislead the classification due to normal clock differences in a cellular network.
In this paper, we propose a WIPS to detect 5G narrowband jammers. This detection is based on spectrograms of the 5G radio signal and on a pre-trained machine-learning model for binary classification. The spectrograms are obtained from an experimental 5G network, which is occasionally attacked by an SSB jammer. The objective of our paper is to compare how accurate and how quickly the SSB jammer can be detected with different machine learning models. To this end, we compare the classification accuracy and computational complexity of a convolutional neural network (CNN) with support vector machines (SVMs) and -nearest neighbors (KNNs). SVM and KNN are combined with principal component analysis (PCA) for dimensionality reduction.
@@
This network is monitored by a watchdog – a separate network element placed within range of the gNB. The watchdog receives radio signals and converts them to the digital baseband in the form of in-phase and quadrature (IQ) samples. On these IQ samples, anomaly detection is performed with a pre-trained machine learning model.
@@
This configuration has several benefits. The watchdog can be deployed independently of the cellular network and, as a mere receiver, does not transmit signals that can be detected. Thus, it can be deployed invisibly to a potential attacker. The watchdog is simpler and easier to construct than a communication device since it can still operate effectively with imperfect synchronization in time and frequency.
II System Assumptions
This section will cover the main assumptions, the experimental setup, and the creation of the dataset.
II-A Main assumptions
We focus on a single cell in a mobile network, including the nodes described above. The gNB, one or multiple UEs, a watchdog, and a jammer are all in range of each other. We assume that the watchdog knows the basic radio parameters such as center frequency, bandwidth, and the pilot structure. This assumption is feasible since the required information is constant and known to the operator of a 5G network.
For each transmitter-receiver path, the baseband-equivalent received signal per subcarrier can be defined as
where is the transmitted orthogonal frequency-division multiplexing (OFDM) symbol, is the gain for the channel between the receiver and the transmitter, and is the additive white Gaussian noise (AWGN) at the receiver. All these quantities are complex signals.
If a jammer starts injecting noise over the same subcarrier, the received signal will become
where is the signal transmitted by the jammer and is the gain for the channel between the receiver and the jammer.
For each subcarrier , the watchdog collects an IQ sample, creating a continuous stream of samples from which spectrograms will be created and analyzed to detect the presence of a jammer.
II-B Studied Jammer
We focus on detecting a narrowband jammer that targets the SSB. Unlike traditional jammers, an SSB jammer does not have to interfere with the complete frequency band in order to disrupt communication. Instead, it generates interference only on those subcarriers used for synchronization. Since , an SSB jammer uses substantially lower bandwidth and energy than traditional jamming. This improves not only the runtime of battery-powered jammers but also heavily reduces their cost, which is mainly defined by Radio Frequency (RF) bandwidth. The comparably low signal bandwidth also makes SSB jamming more difficult to detect than traditional (full band) jamming. These properties make the SSB jammer a serious threat to 5G networks.
As an example, Fig. 1 shows a measured spectrogram of a 5G new radio (NR) signal at center frequency that is jammed on its SSB. Here, the SSB jammer permanently transmits uniform noise at center frequency . The bandwidth of this interference is substantially below the bandwidth of the 5G signal. Note that this signal immediately disappears once the SSB jammer is activated, indicating an effective attack.
II-C Experimental Setup
Fig. 2 shows the experimental setup. We are running a private 5G network in the n78 frequency band, at center frequency MHz, with kHz subcarrier spacing, and multiple-input and multiple-output (MIMO). The system operates in time division duplexing (TDD) mode at a signal bandwidth of 100 MHz or, equivalently, with subcarriers. The SSB covers MHz of bandwidth or, equivalently, uses subcarriers.
The gNB implements the 5G NR air interface as a Software Defined Radio (SDR) by the software srsRAN 23.10 [10] on a personal computer (PC), connected to the universal software radio peripheral (USRP) n300 RF frontend [11]. The transmission power of the gNB is 6 dBm, measured as equivalent isotropic radiated power (EIRP). The UE is a commercial-off-the-shelf 5G modem, namely Quectel RM520N-GL [12]. Its maximum transmission power is dBm (EIRP). Note that the UE adapts its transmission power, while gNB and jammer transmit at constant power.
The core network functionality is provided by the software Open5GS 2.6.6 [13], running on the same PC as srsRAN. This setting provides a 5G standalone network and complies with Release 17 of the 5G standard series 38 [14].
The jammer and the watchdog run on separate computers, each using one USRP X310 [15] RF frontend. The jammer transmits uniform noise with a duty cycle of % at center frequency MHz with MHz of bandwidth and a transmission power of dBm. The watchdog permanently records IQ samples over MHz bandwidth, including a MHz portion of the upper and lower neighboring band, respectively.
The distance between adjacent devices was m. Due to this low value, distance is not a relevant factor in our experiments. Without jamming, the above 5G setup reaches an average throughput of 223 Mbit/s over the air at the transport layer (User Datagram Protocol, UDP). With SSB jamming, neither downlink nor uplink communication is possible.
Dataset creation, training, and measuring the classification time was performed on a single workstation with an Intel Xeon w7-2495X CPU and an NVIDIA RTX A6000 GPU. The operation system was Ubuntu 22.04 LTS with Python 3.9.12 and keras 2.12.0. The GPU was only used for training and is, thus, not utilized during the measurement of classification time.
II-D Dataset Creation
A spectrogram is composed of a stack of arrays, where each array represents the signal’s power spectral density (PSD). Each PSD array is obtained by applying an FFT to a window of IQ samples. We used an FFT rather than Welch’s method [16], sacrificing precision for computational speed with an average error of for each frequency bin of the PDS array. A spectrogram is then composed as a matrix.
Based on previous measurement campaigns, we chose a window size of , leading to a frequency resolution of kHz and a stack of PSDs per spectrogram. This corresponds to a time window of ms.
The resulting spectrogram matrices presented two problems when fed into the training set of the model. First, the power of received radio signals is very low in the linear domain and, thus, typically expressed in the logarithmic domain (decibel). Similarly, we apply the monotonic function to each value of the PSD array at each frequency, to avoid the vanishing gradient problem [17]. Second, due to approximation errors with very low energy bins, some values of the PSDs turned out to be , which causes computation errors with the log function. Then used a small constant , leading to the applied transformation . To each sample , a label is assigned.
II-E Cases and Labeling
@@
The first two cases are classified as legitimate and labeled with , while the third case is classified as an attack and labeled with .
The dataset was then uploaded in IEEE DataPort and is available at[18].
@@ III-A Convolutional Neural Network
As spectrograms are two-dimensional, the first model of choice is a CNN whose structure is given in Table I. The objective is to return the correct label associated to each sample at the final layer of the network which consists in one neuron. The chosen loss function is the binary cross-entropy
@@
Defining as the class of no-jamming and as the class of jamming, the jammer detection is performed by test function
on the input sample with a chosen threshold . We, thus, measure accuracy as probability of false alarm (FA) and misdetection (MD) defined as
@@ III-B Principal Component Analysis
Using spectrograms as model input leads to dimensions. To avoid the curse of dimensionality, we reduce the number of coordinates in two steps.
@@
Second, we further reduce the dimensions with principal component analysis (PCA). The basic idea of PCA is to change the basis so that the first coordinates (or principal components) contain “the most information” about the underlying data. The ratio of explained variance measures the information contained in each component and sums up to when summed over all components. The cumulated ratio of explained variance is displayed in Fig. 3. It can be seen that a few components already capture most of the information. In particular, 98%, 98.5%, and 99% of the variance are explained by , , and principal components, respectively. This suggests components an optimal trade-off between computational complexity and accuracy.
@@ III-D Support Vector Machines
SVMs are a powerful class of machine learning models, which can be used for supervised classification and regression. SVMs operate by constructing a decision hyperplane in the feature space that optimally separates different classes. By employing the kernel trick, support vector machines can also be used effectively for non-linear tasks. In this work, we employed SVMs with various kernels after dimension reduction with PCA.
@@ IV-A Convolutional Neural Network
The training set was composed of samples, equally distributed among the three cases described in Sec. II-E. Using the same distribution, the validation set was composed of samples. This set is used to monitor the validation loss and to stop the training phase after three increases of the parameter. The test set contains samples, distributed as the training and validation set.
@@
To study the performance in a wider range of scenarios while still maintaining experimental control, we reduced the gain of the jammer from 80 dB to 45 dB in several steps. Assuming a dominant direct path with distance and quadratic path loss , this corresponds to an increase of the distance between jammer and watchdog from 1 to 56.23 m. Note that a gain of 80 dB corresponds to the standard scenario and that all other parameters remain as in Section II-C.
Fig. 6 shows the resulting accuracy for these experiments as the 90th percentile of the CNN output with an active jammer. In this situation, the correct output of the CNN would be 1 while 0 would be a false negative and no other output values are allowed. Consequently, the closer the shown percentile is to 1, the higher the accuracy. The results show such high accuracy for gains of 65 dB and above. With above path loss model, this means that jammers within m distance from the watchdog’s antenna can be detected at high accuracy. For larger distances, the accuracy quickly drops to a false negative rate of 100%.
Note that, due to the low transmission power of our laboratory equipment, the absolute distance value is not relevant in such geometry discussions. Instead, the ratios between gains (or, equivalently, distances) need to be considered. This means that jammers can still be accurately detected if they exceed the UE-to-basestation distance by a factor of . This is an impressive result, although the sudden accuracy drop points to an over-fitted model which can be improved by retraining with various jamming distances.
@@
For each of the training, validation, and test sets, 1024 samples per class were selected from each of the three cases in Sec. II-E. The training, validation, and test accuracy for the settings (i) standardized PSDs and (ii) projections of non-standardized PSDs onto the first 8 principal components are displayed in Table II. The optimal choice of neighbors based on the validation data is for setting (i) and for setting (ii). With an accuracy of approximately %, we can reliably detect jammers.
Fig. 7 shows the CDF of the classification time of the KNN. With standardization, the model classified of the cases within ms. With PCA, the model performed substantially better with a classification time of ms for of the cases.
@@ IV-C Support Vector Machines
Our experiments with SVMs showed that this class of models is well-suited for jamming detection. The overall accuracy on the validation and test dataset significantly depends on the used kernel. We considered linear, polynomial, and radial basis function (RBF) kernels. On the other hand, increasing the number of dimensions does not show a significant effect.
@@
This large dimension reduction highly reduces the computational complexity of the classification. Fig. 8 shows that the SVM model classified 95% of the cases within ms. Thus, SVMs can detect jammers using substantially smaller computation time than the CNN (Fig. 5) and than the KNN without PCA (Fig. 7).
@@ V Conclusion
All three models yielded very high accuracy. CNN correctly classified % of the jammed cases, followed by test scores of up to % with SVM, and by test scores of up to % with KNN. Thus, in terms of accuracy, either of these methods can be chosen.
Particularly interesting are the results for computation time due to their substantial differences between the machine learning methods. While the CNN required 38 ms to classify 95% of the cases, KNN needed only 8 ms if combined with PCA. The fasted classification was achieved by SVM with PCA, with a classification time of ms for 95% of the cases.
Of further interest is a first impression on detection performance for variable distances. While being limited to the CNN, the results already show that the watchdog does not have to be close to the jammer in order to detect it. The sudden drop for large distances points to an over-fitted model. This, however, can be improved by retraining with various jamming distances.
Motivated by these promising results, we will now study the robustness of the detection accuracy for various wireless network geometries and for multiple jamming signals.